OpenStudy (anonymous):

10^-1/2 instructions would help

OpenStudy (anonymous):

Firstly, it's negative so that means you write it as the reciprocal with a positive exponent: \[\frac{ 1 }{ 10^{1/2} }\] You can convert fractional powers to roots by doing this: the numerator of the fraction becomes the exponent and the denominator of the fraction the "nth" root. So your answer is \[\frac{ 1 }{ \sqrt{10} }\]

OpenStudy (anonymous):

thanks!